# Position: Foundation Agents as the Paradigm Shift for Decision Making

### v1

@@ Abstract
Decision making demands intricate interplay between perception, memory, and reasoning to discern optimal policies. Conventional approaches to decision making face challenges related to low sample efficiency and poor generalization. In contrast, foundation models in language and vision has showcased rapid adaptation to diverse new tasks. Therefore, we advocate for the construction of foundation agents as a transformative shift in the learning paradigm of agents. This proposal is underpinned by the formulation of foundation agents with its fundamental characteristics and challenges motivated by the success of large language models (LLMs). Moreover, we specify the roadmap of foundation agents from large interactive data collection or generation, to self-supervised pretraining and adaptation, and knowledge and value alignment with LLMs. Lastly, we pinpoint critical research questions derived from the formulation and delineate trends for foundation agents supported by real-world use cases, addressing both technical and theoretical aspects to propel the field towards a more comprehensive and impactful future.
@@ 1 Introduction
To arrive at this position, we first identify three fundamental characteristics of foundation agents: (1) a unified representation of variables involved in decision making, including state-action spaces, feedback signals (e.g., rewards or goals) and environment dynamics, (2) a unified policy interface across tasks and domains from robotics and game play to healthcare and beyond, and (3) interactive decision-making in physical and virtual worlds by reasoning about behaviors, handling environment stochasticity and uncertainty, and potentially navigating competitive or cooperative multi-agent scenarios. These characteristics constitute the uniqueness and challenges for foundation agents [^1] Distinctions from task-specific agents and recent LLM-based agents can be found in Appendix A., empowering them with multi-modality perception, multi-task and cross-domain adaptation as well as few- or zero-shot generalization. Particularly, foundation agents exhibit enhanced credit assignment and planning in scenarios requiring long-horizon reasoning (Wei et al., 2022; Yao et al., 2022, 2023) and involving sparse rewards or partial observability (Meng et al., 2021; Guo et al., 2023). Such unique set of characteristics and capabilities enable foundation agents to improve sample efficiency and generalization, showing their versatility in navigating diverse contexts of decision making.
@@ 2.2 Self-Supervised Learning for RL
To improve sample efficiency and generalization, conventional self-supervised learning for RL involves representation learning of separate RL components, such as action representation (Gu et al., 2022), reward representation (Ma et al., 2022), policy representation (Tang et al., 2022), and environment or task representation (Sang et al., 2022). Recently, by formulating RL as a sequence modeling problem (Chen et al., 2021; Janner et al., 2021), the representation of various RL components can be simultaneously learned via trajectory optimization (Lee et al., 2022; Liu et al., 2022; Carroll et al., 2022; Sun et al., 2023; Wu et al., 2023; Boige et al., 2023). Such a formulation draws upon the simplicity and scalability of Transformer (Vaswani et al., 2017), and benefits representation learning of knowledge concerning different aspects of sequential decision-making processes.
@@ 3 Learning from Large-Scale Interactive Data
Large-scale interactive data is an essential component for building foundation agents, akin to the significance of internet text and image data for foundation models in language and vision. In this section, we first demonstrate how a potential foundation agent can be trained when large, multi-modal and multi-task demonstrations are available. We then discuss the potential use of data generation systems or real-world simulators for training foundation agents. Finally, we identify the constraints and alternatives to offline RL or IL methods in establishing foundation agents, as well as the limitations of real-world simulators realized by video generation models.
@@ 3.2 Potentials of Real-World Simulators
Generative models as data generation systems (Mandlekar et al., 2023) or real-world simulators (Yang et al., 2023a; OpenAI, 2024; Bruce et al., 2024) open another path for training foundation agents through imitating behaviors from generated data. For instance, RoboGen (Wang et al., 2023c) realizes a propose-generate-learn cycle to scale up robot learning via generative simulation, of which policy learning algorithms are not restricted to IL but also involve motion planning or trajectory optimization. Moreover, video generation models such as Sora (OpenAI, 2024) and Genie (Bruce et al., 2024) can be a general-purpose simulator of interactive decision-making scenarios in physical and virtual worlds. These models, by unifying the representation of visual worlds (Yang et al., 2024), generate diverse and open-ended training environments across tasks and domains for scaling up the foundation agent in a never-ending curriculum of new, generated worlds. The foundation agent pretrained with these generated environments learns to reason about behaviors, handle environment stochasticity and uncertainty, and navigate competitive or cooperative multi-agent settings, thus enhancing adaptability and generalization in unseen scenarios. This could be particularly valuable for domains where real-world interactive data is scarce or expensive to obtain, such as robotics and self-driving.
Table 1: Self-supervised pretraining objectives for decision making with Transformer. $P_{\theta}$ is the prediction of the pretrained model. $\tau_{0:t-1}$ denotes previous trajectories before timestep $t$. $s_{t}$ is the state or observation at timestep $t$ and $a_{t}$ the action at timestep $t$. $R_{t}$ is the return (cumulative rewards) at timestep $t$, while $\hat{R_{t}}$ is the predicted return and $r_{t}$ the predicted immediate reward at timestep $t$. $G_{t+i}$ refers to a sub-goal or the ultimate goal at future timesteps, and $z$ denotes the encoded future trajectory in the same sequence length as the input trajectory. Note that the conditional variables in each loss function are encoded by either a causal Transformer (Brown et al., 2020) for autoregressive modeling (top rows) or a bidirectional Transformer (Devlin et al., 2018) for masked modeling (bottom rows).
@@ 4.1 Self-Supervised Pretraining
Self-supervised (unsupervised) pretraining for sequential decision making allows foundation agents to learn without reward signals and encourages the agent to learn from suboptimal offline datasets. This is particularly applicable when large, unlabeled data can be easily collected from internet or real-world simulators. Specifically, given a sequence of trajectories $\tau=\{(s_{1},a_{1},s_{2},a_{2},,...,s_{T},a_{T})\}^{N}$, self-supervised pretraining aims to learn a representation function $g:\mathbb{T}\in\mathbb{R}^{d}\to\mathbb{Z}\in\mathbb{R}^{m}$ $(m\ll d)$ to distill valuable knowledge from trajectory data for downstream inference. The knowledge can be temporal information about the same modality (e.g., $s_{t}\rightarrow s_{t+i}$), causal information between different modalities (e.g., $\pi(a|s)$), as well as dynamics $P(s^{\prime}|s,a)$ and reward $R(s,a)$ information.
@@
The second step is devising self-supervised pretraining objectives to discover the underlying structure and semantics of trajectory data for knowledge representation learning. Table 1 shows pretraining objectives for sequential decision making in domains including control (Sun et al., 2023), navigation (Carroll et al., 2022), and game play (Lee et al., 2022). These objectives are mainly inspired by autoregressive (Radford et al., 2019; Brown et al., 2020) and masked (Devlin et al., 2018; He et al., 2022) prediction in language and vision model pretraining. The primary pretraining objective for decision making involves learning control information by predicting the next action in an autoregressive way. This objective is further modified by conditioning on different variables within trajectories, such as reward or value signals (Chen et al., 2021; Lee et al., 2022), next state (observation) information (Sun et al., 2023), or latent future sub-trajectory information (Xie et al., 2023). As an alternative, random masking (Liu et al., 2022) learns the context of trajectory data by filling in missing information for pretraining. Various masking schemes conditioned on different RL components are designed. For example, reward-conditioned random mask prediction (Carroll et al., 2022) recovers masked trajectory segments conditioned on the first-step return, while (Sun et al., 2023) only recovers masked actions conditioned on the final-step action to capture global temporal relations for multi-step control. Combining masked and autoregressive modeling, (Wu et al., 2023) constrains the last variable in trajectories to be masked to force the pretrained model to be causal at inference time. Moreover, contrastive prediction objective (Stooke et al., 2021; Schwarzer et al., 2021; Yang & Nachum, 2021; Cai et al., 2023) has been commonly used in self-supervised pretraining to learn state representation via a contrastive loss, which can benefit dynamics learning.
@@ 4.2 Downstream Adaptation
In adaptation, extensive decision-related knowledge acquired from pretraining is transferred to downstream tasks via fine-tuning or prompting. The knowledge transfer facilitates the optimization of learning objectives $f_{\theta}(z)$ in downstream inference, such as the value function $V(\pi)$ or $Q(s,a)$, policy, dynamics and reward functions. The optimized objectives empowered by knowledge reasoning can finally improve sample efficiency and generalization compared to learning from scratch in traditional RL.
@@ 5.1 Memory and Information Processing
Foundation agents aligned with LLMs is often equipped with a memory module. The memory module functions as a repository, storing both task-specific information and past interaction history that enables agents to retrieve and refine relevant historical data for future planning. Similar as human memory (Izquierdo et al., 1999), the agent’s memory can be categorized into short-term and long-term components. Short-term memory encapsulates recent information within the context window, while long-term memory encompasses historical information stored in external storage.
@@ 5.3 Action and Decision Making
For ultimate action execution and interaction with real or simulated environments, an action module is required for foundation agents to interpret plans from the planning module, translates subgoals into executable actions, and combine atomic actions to form more complex structured actions.
@@ 6.1 A Unified or Compositional Foundation Agent
One central challenge in building foundation agents stems from the substantial diversity among sequential decision making tasks, involving variations in state and action spaces, environment dynamics, and reward functions. In contrast, the unification of tasks in language is facilitated by the text-to-text generation paradigm (Brown et al., 2020; OpenAI, 2023). To structure diverse environments and tasks, some attempts have been made along two primary directions: (1) the pursuit of a unified foundation model (Reed et al., 2022; Brohan et al., 2022), and (2) the exploration of compositions involving multiple existing foundation models (Zeng et al., 2022; Ajay et al., 2023) for decision making.
@@ 6.2 Policy Optimization with Foundation Agents
Establishing theoretical foundations for policy optimization with foundation agents is a complex but crucial endeavor. Firstly, policies generated by foundation agents will directly affect changes in real-world applications, requiring rigorous theoretical guarantees to ensure their effectiveness, optimality, safety and robustness. Meanwhile, the theoretical foundations for policy optimization can in turn help to understand the principles underlying the agent’s behavior and performance, thereby enabling the development of algorithms and methodologies for training foundation agents. Specifically, we posit that the theoretical foundations require an interdisciplinary view (e.g., control theory, RL, optimization.), and we outline some ideas here.
@@
  • Understand the Interplay between Pretraining and Task-Specific Optimization: Establish a mathematical framework to investigate how the pretraining phase shapes the optimization landscape for task-specific policy optimization and how different adaptation algorithms interact with the previously acquired knowledge. This could involve probing convergence properties, sample efficiency, and generalization bounds of foundation agents.
  • Extend Existing Theoretical Frameworks: Adapt and extend existing theoretical frameworks from RL, control theory and optimization to accommodate the peculiarities of foundation agents. This may involve relaxing assumptions made in traditional RL and developing new notions of optimality for generalist policies. For example, (1) extend the MDP Framework with generalized state spaces to incorporate multi-modal inputs and extended action spaces spanning different domains (e.g., physical movements, linguistic responses), (2) leverage the probabilistic representations and principled reasoning offered by the control as inference framework (Todorov, 2008) to derive optimal control policies and analyze their properties, and (3) ground policy optimization for foundation agents in information-theoretic principles, such as mutual information or entropy, to capture the agent’s ability to influence the environment and gather informative experiences.
@@ 6.3 Open-Ended Tasks for Foundation Agents
In the evolution of foundation agents, a notable trend is the shift towards learning from open-ended tasks at scale (Team et al., 2021; Fan et al., 2022). Traditionally, RL agents are confined to a single, individual task (Silver et al., 2017; Schwarzer et al., 2023), limiting their applicability to massive dynamic scenarios. Contemporary approaches instead prioritize the scalability of generative models, especially via LLMs to effectively handle the complexities inherent in open-ended tasks (Wang et al., 2023a; Zhu et al., 2023). The open-ended agents, endowed with continuously learning capabilities, interact with environments or external tools (Shen et al., 2023; Hu et al., 2023; Ge et al., 2023), iteratively refining their decision making process. This paradigm aligns seamlessly with RL principles, where agents acquire optimal strategies through iterative learning.
@@
  • Creativity and Innovative Solutions: Generating novel and valuable solutions is a hallmark of open-ended learning. This necessitates agents that can go beyond mere imitation or optimization, possibly by combining known concepts in unique ways or inventing entirely new ones. Potential solutions include (1) Quality-Diversity algorithms (Grillotti & Cully, 2022) to produce a diverse set of high-performing solutions, (2) neuro-evolutionary methods (Lehman & Stanley, 2011) to produce innovative solutions, and (3) deep generative models to generate novel and sometimes unexpectedly creative outputs.
@@ 7 Conclusion
We take the position that foundation agents hold the potential to alter the landscape of agent learning for decision making, akin to the revolutionary impact of foundation models in language and vision. We support this position by formulating the notion of foundations agents, specify their roadmap and key challenges. Demonstrated by recent work and a case study, we show that the enhanced perception, adaptation, and reasoning abilities of agents not only address limitations of conventional RL, but also hold the key to unleash the full potential of foundation agents in real-world decision making. However, due to the intrinsic complexity of decision making, existing work towards foundation agents are still at a starting point. Future directions involve but not limited to learning a unified representation of diverse environments and tasks, establishing theoretical foundations for a unified policy interface and learning from open-ended tasks at scale. Bridging these gaps will pave the way for more robust and versatile foundation agents, contributing to the evolution of artificial general intelligence.

### v2

@@ Abstract
Decision making demands intricate interplay between perception, memory, and reasoning to discern optimal policies. Conventional approaches to decision making face challenges related to low sample efficiency and poor generalization. In contrast, foundation models in language and vision have showcased rapid adaptation to diverse new tasks. Therefore, we advocate for the construction of foundation agents as a transformative shift in the learning paradigm of agents. This proposal is underpinned by the formulation of foundation agents with their fundamental characteristics and challenges motivated by the success of large language models (LLMs). Moreover, we specify the roadmap of foundation agents from large interactive data collection or generation, to self-supervised pretraining and adaptation, and knowledge and value alignment with LLMs. Lastly, we pinpoint critical research questions derived from the formulation and delineate trends for foundation agents supported by real-world use cases, addressing both technical and theoretical aspects to propel the field towards a more comprehensive and impactful future.
@@ 1 Introduction
To arrive at this position, we first identify three fundamental characteristics of foundation agents: (1) a unified representation of variables involved in decision process, including state-action spaces, feedback signals (e.g., rewards or goals) and environment dynamics, (2) a unified policy interface across tasks and domains from robotics and game play to healthcare and beyond, and (3) interactive decision-making in physical and virtual worlds by reasoning about behaviors, handling environment stochasticity and uncertainty, and potentially navigating competitive or cooperative multi-agent scenarios. These characteristics constitute the uniqueness and challenges for foundation agents [^1] Distinctions from task-specific agents and recent LLM-based agents can be found in Appendix A., empowering them with multi-modality perception, multi-task and cross-domain adaptation as well as few- or zero-shot generalization. Particularly, foundation agents exhibit enhanced credit assignment and planning in scenarios requiring long-horizon reasoning (Wei et al., 2022; Yao et al., 2022, 2023) and involving sparse rewards or partial observability (Meng et al., 2021; Guo et al., 2023). Such unique set of characteristics and capabilities enable foundation agents to improve sample efficiency and generalization, showing their versatility in navigating diverse contexts of decision making.
@@ 2.2 Self-Supervised Learning for RL
To improve sample efficiency and generalization, conventional self-supervised learning for RL involves representation learning of separate RL components, such as action representation (Gu et al., 2022), reward representation (Ma et al., 2022), policy representation (Tang et al., 2022), and environment or task representation (Sang et al., 2022). Recently, by formulating RL as a sequence modeling problem (Chen et al., 2021; Janner et al., 2021), the representation of various RL components can be simultaneously learned via trajectory optimization (Lee et al., 2022; Liu et al., 2022; Carroll et al., 2022; Sun et al., 2023; Wu et al., 2023; Boige et al., 2023). Such a formulation draws upon the simplicity and scalability of Transformer (Vaswani et al., 2017), and benefits representation learning of knowledge concerning different aspects of decision process.
@@ 3 Learning from Large-Scale Interactive Data
Large-scale interactive data is an essential component for building foundation agents, akin to the significance of internet text and image data for foundation models in language and vision. In this section, we first demonstrate how a potential foundation agent can be trained via offline IL when large, multi-modal and multi-task demonstrations are available. We then discuss the potential use of data generation systems or real-world simulators for training foundation agents at scale. Finally, we identify the constraints and alternatives to offline RL or IL methods in establishing foundation agents, as well as the limitations of real-world simulators realized by video generation models.
@@ 3.2 Potentials of Real-World Simulators
Generative models as data generation systems (Mandlekar et al., 2023) or real-world simulators (Yang et al., 2023a; OpenAI, 2024; Bruce et al., 2024) open another path for training foundation agents through learning behaviors from generated data. For instance, RoboGen (Wang et al., 2023c) realizes a propose-generate-learn cycle to scale up robot learning via generative simulation, of which policy learning algorithms are not restricted to IL but also involve motion planning or trajectory optimization. Moreover, video generation models such as Sora (OpenAI, 2024) and Genie (Bruce et al., 2024) can be a general-purpose simulator of interactive decision making scenarios in physical and virtual worlds. These models, by unifying the representation of visual worlds (Yang et al., 2024), generate diverse and open-ended training environments across tasks and domains for scaling up the foundation agent in a never-ending curriculum of new and generated worlds. The foundation agent pretrained with these generated environments learns to reason about behaviors, handle environment stochasticity and uncertainty, and navigate competitive or cooperative multi-agent settings, thus enhancing adaptability and generalization in unseen scenarios. This could be particularly valuable for domains where real-world interactive data is scarce or expensive to obtain, such as robotics and self-driving.
Table 1: Self-supervised pretraining objectives for decision making with Transformer. $P_{\theta}$ is the prediction of the pretrained model. $\tau_{0:t-1}$ denotes previous trajectories before timestep $t$. $s_{t}$ is the state or observation at timestep $t$ and $a_{t}$ the action at timestep $t$. $R_{t}$ is the return-to-go (target return for the rest of the input sequence) at timestep $t$, while $\hat{R_{t}}$ is the predicted future value and $r_{t}$ the future reward. $G_{t+i}$ refers to a sub-goal or the ultimate goal at future timesteps, and $z$ denotes the encoded future trajectory in the same sequence length as the input trajectory. Note that the conditional variables in each loss function are encoded by either a causal Transformer (Brown et al., 2020) for autoregressive modeling (top rows) or a bidirectional Transformer (Devlin et al., 2018) for masked modeling (bottom rows).
@@ 4.1 Self-Supervised Pretraining
Self-supervised (unsupervised) pretraining for decision making allows foundation agents to learn without reward signals and encourages the agent to learn from suboptimal offline datasets. This is particularly applicable when large, unlabeled data can be easily collected from internet or real-world simulators. Specifically, given a sequence of trajectories $\tau=\{(s_{1},a_{1},s_{2},a_{2},,...,s_{T},a_{T})\}^{N}$, self-supervised pretraining aims to learn a representation function $g:\mathbb{T}\in\mathbb{R}^{d}\to\mathbb{Z}\in\mathbb{R}^{m}$ $(m\ll d)$ to distill valuable knowledge from trajectory data for downstream inference. The knowledge can be temporal information about the same modality (e.g., $s_{t}\rightarrow s_{t+i}$), causal information between different modalities (e.g., $\pi(a|s)$), as well as dynamics $P(s^{\prime}|s,a)$ and reward $R(s,a)$ information.
@@
The second step is devising self-supervised pretraining objectives to discover the underlying structure and semantics of trajectory data for knowledge representation learning. Table 1 shows pretraining objectives for decision making in domains including control (Sun et al., 2023), navigation (Carroll et al., 2022), and game play (Lee et al., 2022). These objectives are mainly inspired by autoregressive (Radford et al., 2019; Brown et al., 2020) or masked (Devlin et al., 2018; He et al., 2022) prediction in language and vision model pretraining. The primary pretraining objective for decision making involves learning control information by predicting the next action in an autoregressive way. This objective is further modified by conditioning on different variables within trajectories, such as reward or value signals (Chen et al., 2021; Lee et al., 2022), next state (observation) information (Sun et al., 2023), or latent future sub-trajectory information (Xie et al., 2023). As an alternative, random masking (Liu et al., 2022) learns the context of trajectory data by filling in missing information. Various masking schemes conditioned on different RL components are designed. For example, reward-conditioned random mask prediction (Carroll et al., 2022) recovers masked trajectory segments conditioned on the first-step return, while (Sun et al., 2023) only recovers masked actions conditioned on the final-step action to capture global temporal relations for multi-step control. Combining masked and autoregressive modeling, (Wu et al., 2023) constrains the last variable in trajectories to be masked to force the pretrained model to be causal at inference time. Moreover, contrastive prediction objective (Stooke et al., 2021; Schwarzer et al., 2021; Yang & Nachum, 2021; Cai et al., 2023) has been commonly used in self-supervised pretraining to learn state representation via a contrastive loss, which can benefit dynamics learning.
@@ 4.2 Downstream Adaptation
During adaptation, extensive decision-related knowledge acquired from pretraining is transferred to downstream tasks via fine-tuning or prompting. The knowledge transfer facilitates the optimization of learning objectives $f_{\theta}(z)$ in downstream inference, such as the value function $V(\pi)$ or $Q(s,a)$, policy, dynamics and reward functions. The optimized objectives empowered by knowledge reasoning can finally improve sample efficiency and generalization compared to learning from scratch in traditional RL.
@@ 5.1 Memory and Information Processing
Foundation agents aligned with LLMs are often equipped with a memory module. The memory module functions as a repository, storing both task-specific information and past interaction history that enables agents to retrieve and refine relevant historical data for future planning. Similar as human memory (Izquierdo et al., 1999), the agent’s memory can be categorized into short-term and long-term components. Short-term memory encapsulates recent information within the context window, while long-term memory encompasses historical information stored in external storage.
@@ 5.3 Action and Decision Making
For ultimate action execution and interaction with real or simulated environments, an action module is required for foundation agents to interpret plans from the planning module, translate subgoals into executable actions, and combine atomic actions to form more complex structured actions.
@@ 6.1 A Unified or Compositional Foundation Agent
One central challenge in building foundation agents stems from the substantial diversity among decision making tasks, involving variations in state and action spaces, environment dynamics, and reward functions. In contrast, the unification of tasks in language is facilitated by the text-to-text generation paradigm (Brown et al., 2020; OpenAI, 2023). To structure diverse environments and tasks, some attempts have been made along two primary directions: (1) the pursuit of a unified foundation model (Reed et al., 2022; Brohan et al., 2022), and (2) the exploration of compositions involving multiple existing foundation models (Zeng et al., 2022; Ajay et al., 2023) for decision making.
@@ 6.2 Policy Optimization with Foundation Agents
Establishing theoretical foundations for policy optimization with foundation agents is a complex but crucial endeavor. Firstly, policies generated by foundation agents will directly affect changes in real-world applications, requiring rigorous theoretical guarantees to ensure their effectiveness, optimality, safety and robustness. Meanwhile, the theoretical foundations for policy optimization can in turn help to understand the principles underlying the agent’s behavior and performance, thereby enabling the development of algorithms and methodologies for training foundation agents. Specifically, we posit that the theoretical foundations require an interdisciplinary view (e.g., control theory, RL, and optimization), and we outline some ideas here.
@@
  • Understand the Interplay between Pretraining and Task-Specific Optimization: Establish a mathematical framework to investigate how the pretraining phase shapes the optimization landscape for task-specific policy optimization, and how different adaptation algorithms interact with the previously acquired knowledge. This could involve probing convergence properties, sample efficiency, and generalization bounds of foundation agents.
  • Extend Existing Theoretical Frameworks: Adapt and extend existing theoretical frameworks from RL, control theory and optimization to accommodate the peculiarities of foundation agents. This may involve relaxing assumptions made in traditional RL and developing new notions of optimality for generalist policies. For example, (1) extend the MDP framework with generalized state spaces to incorporate multi-modal inputs and extended action spaces spanning different domains (e.g., physical movements and linguistic responses), (2) leverage the probabilistic representations and principled reasoning offered by the control as inference framework (Todorov, 2008) to derive optimal control policies and analyze their properties, and (3) ground policy optimization for foundation agents in information-theoretic principles, such as mutual information or entropy, to capture the agent’s ability to influence the environment and gather informative experiences.
@@ 6.3 Open-Ended Tasks for Foundation Agents
In the evolution of foundation agents, a notable trend is the shift towards learning from open-ended tasks at scale (Team et al., 2021; Fan et al., 2022). Traditionally, RL agents are confined to a single and individual task (Silver et al., 2017; Schwarzer et al., 2023), limiting their applicability to massive dynamic scenarios. Contemporary approaches instead prioritize the scalability of generative models, especially via LLMs to effectively handle the complexities inherent in open-ended tasks (Wang et al., 2023a; Zhu et al., 2023). The open-ended agents, endowed with continuously learning capabilities, interact with environments or external tools (Shen et al., 2023; Hu et al., 2023; Ge et al., 2023), iteratively refining their decision process. This paradigm aligns seamlessly with RL principles, where agents acquire optimal strategies through iterative learning.
@@
  • Creativity and Innovative Solutions: Generating novel and valuable solutions is a hallmark of open-ended learning. This necessitates agents that can go beyond mere imitation or optimization, possibly by combining known concepts in unique ways or inventing entirely new ones. Potential solutions include (1) Quality-Diversity algorithms (Grillotti & Cully, 2022) to produce a diverse set of high-performing solutions, (2) neuroevolution methods (Lehman & Stanley, 2011) to produce innovative solutions, and (3) deep generative models to generate novel and sometimes unexpectedly creative outputs.
@@ 7 Conclusion
We take the position that foundation agents hold the potential to alter the landscape of agent learning for decision making, akin to the revolutionary impact of foundation models in language and vision. We support this position by formulating the notion of foundations agents and specifying their roadmap and key challenges. Demonstrated by recent work and a case study, we show that the enhanced perception, adaptation, and reasoning abilities of agents not only address limitations of conventional RL, but also hold the key to unleash the full potential of foundation agents in real-world decision making. However, due to the intrinsic complexity of decision making, existing work towards foundation agents is still at a starting point. Future directions involve but are not limited to learning a unified representation of diverse environments and tasks, establishing theoretical foundations for a unified policy interface and learning from open-ended tasks at scale. Bridging these gaps will pave the way for more robust and versatile foundation agents, contributing to the evolution of artificial general intelligence.
Acknowledgements
This work is supported in part by the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No.XDA27010103) and the Youth Innovation Promotion Association CAS.